Question:
x(squared)-14x+24=0

Answer by tutorcecilia(2152)

x^2-14x+24=0
(x-2)(x-12)=0 [factor]
x-2=0 [set each factor equal to zero]
x=2 [solve for x]
.
x-12=0
x=12
.
Check by plugging either x=2 or x=12 back into the orginal equation.